solve the system of equation
x+y+7z=-7
x+y+2z=-10
x-8y-3z=50

Answer by juliane1(5)

1. x + y + 7z = -7
(x + y + 7z) + (-y-7z)
= -7+(-y - 7z)
x + y + 7z - y - 7z - 7 - y - 7z
x + y - y + 7z - 7z = -y-7z-7
x=-y-7z-7
2. x+y+2z=-10
(x+y+2z) + -y-2z)= -10+(-y-2z)
x+y+2z-y-2z=-10-y-2z
x+y-y+2z = -y-2z-10
x = -y-2z-10
3. x-8y-3z=50
(x-8y-3z) + (8y + 3z) = 50 + (8y + 3z)
x - 8y - 3z + 8y + 3z = 50 + 8y + 3z
x - 8y + 8y -3z + 3z + 8y + 3z + 50
8 = 8y + 3z + 50